A person can finish his building work in 8days, but unfortunately while building every day he loses his 20% of work. inthat case how many days he can finish his building work?

Answer by Edwin McCravy(20060)

Each day he does 1/8th of the job but then loses 20% of that 1/8.

Since 20% = 20/100 = 1/5, he loses 1/5th of 1/8th of the job.

1/5th of 1/8th is 1/40th.  So each day he loses 1/40th of the job 

out of that 1/8th that he did.  So each day he completes 

1/8 - 1/40 = 5/40 - 1/40 = 4/40 = 1/10th of the job.

So it will take him 10 days to finish the job.

Edwin